OpenStudy (anonymous):

Gr12 Calculus question, please help.

OpenStudy (anonymous):

A circular oil spill on the surface of the ocean spreads outward. Find the approximate rate of change in the area of the oil slick with respect to its radius when the radius is 100 m.

OpenStudy (freckles):

so we need to know the area of circle formula thingy

OpenStudy (anonymous):

A=2pir?

OpenStudy (freckles):

\[A=\pi r^2 \]

OpenStudy (freckles):

now do you know how to differentiate both sides w.r.t to the variable t. t represents time.

OpenStudy (anonymous):

no, sorry i forgot how to do that :( do we use limits?

OpenStudy (freckles):

pi is a constant multiple so you have \[\frac{dA}{dt}=\pi \frac{dr^2}{dt}\]

OpenStudy (freckles):

to find the derivative of r^2 use chain rule power rule combination

OpenStudy (anonymous):

And i would find my answer after that or would there be a next step? :(

OpenStudy (freckles):

your goal is to find approximate A' when r=100

OpenStudy (anonymous):

The answer on my book says 200pir^2 m^2/m

OpenStudy (anonymous):

How would i find the units when I'm only given m?

OpenStudy (freckles):

that r in the answer makes no sense to me

OpenStudy (freckles):

did you mean 200 pi*m^2/m ?

OpenStudy (anonymous):

Oh yes, no r. Sorry!

OpenStudy (freckles):

I will talk about the units but first can you tell the derivative of r^2?

OpenStudy (freckles):

Ok I guess I will explain the units since you don't need help with the other part. Area is measured with a square unit Length is measured with a non square unit So if your length is measured in m Then your area is measured in m^2 A- > m^2 r-> m and you have to find dA/dr so therefure your units is m^2/m

OpenStudy (freckles):

\[A(r)=\pi r^2 \\ \frac{dA}{dr}= \pi \frac{d(r)^2}{dr}=?\]